OpenStudy (kimberly101):

Goods listed at $180 were discounted by 22%. If they had sold for $100, what would be the percentage discount have been?

OpenStudy (anonymous):

List price is $ 180 and selling price is $ 100 so disc = $ 180- $ 100 =$ 80 So percentage discount. Is.( $ 80/ $ 180 )x 100 = 44.4%

OpenStudy (anonymous):

180(1-22/100)=100(1-x/100) 180(78)=100(100-x) nw u calculate x